Circular No. 8253

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V838 MONOCEROTIS M. Mikolajewski and T. Tomov, Uniwersytet Mikolaja Kopernika; R.

Kurtev, Sofia University; L. Georgiev, Universidad Nacional Autonoma de Mexico; and A. Wolszczan, Pennsylvania State University, write that they obtained UBV I and Hα images of V838 Mon and its light echo with the 1.5m telescope at San Pedro Martir on Oct. 18.5 UT, noting that V838 Mon recently has become bluer (cf. IAUC 8202, 8210) and comparable in bands U and B to its previously reported B3-V companion (IAUC 8005). V838 Mon itself, and an additional star ∼ 1′′ to the southeast, are clearly visible in their U images. The B-band centroid of the V838 Mon image is shifted in the same direction, while frames at longer wavelengths display only one star at this position. The authors’ low-resolution spectral observations of V838 Mon obtained on Jan. 27.2 and Mar. 24.1 with the 10-m Hobby-Eberly Telescope (range 430–730 nm, resolving power 650) also show a composite spectrum. The blue and yellow regions are dominated by a reddened (V ∼ 15.9; B ∼ 16.5), normal B3-V star. The red part of the spectrum is characterized by strong features belonging to an extremely cool (L-type?) supergiant, evidently a remnant of the V838 Mon outburst. The analyses of the interstellar features and reddening suggest that the B3-V star is a blending objects lying ∼ 1000 pc behind V838 Mon. Two rings are apparent in the current light echo visible on the new V, I, and Hα images: an almost concentric outer circle with radius ∼ 50′′ and an eccentric inner ellipse with half-axes of ∼ 25′′ × 32′′.

COMET 158P/KOWAL-LINEAR Following the reported identification on IAUC 8247 of P/2001 RG_{100}

(LINEAR) with C/1979 O1 (Kowal), the IAU Committee on Small-Body Nomenclature has agreed to rename this comet 158P/Kowal-LINEAR.

NOVAE IN M31 Further to IAUC 8238, M. Fiaschi, D. Tiveron, and L. Milani report the

discovery of an apparent nova located at α = 0^{h}43^{m}00^{s}.75, δ = +41^{o}11′26′′.7 (equinox 2000.0), which is 185′′ east and 282′′ east of the center of M31. Available Hα magnitudes: Nov. 29, [18; Dec. 3.722 UT, 16.6; 5.776, 15.8. They also report Hα magnitudes of the new apparent nova reported on IAUC 8248: Nov. 4.966, 17.2; 29.760, 17.0; Dec. 3.722, 17.3; 5.776, 18:.

2003 December 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green